31) If tan theta= 0.75, and 0 to 90 deg, evaluate sin theta; and cos theta.
Answer by stanbon(75887)

If tan theta= 0.75, and 0 to 90 deg, evaluate sin theta; and cos theta.
------
tan(x) = y/x = 3/4 in QI where x and y are both positive.
So y = 3 and x = 4
Therefor r = sqrt[3^2+4^2] = 5
-----
sin = y/r = 3/5
cos = x/r = 4/5
tan = y/x = 3/4
---
csc = 5/3
sec = 5/4
cot = 4/3
===============
Cheers,
Stan H.
=====================
---